# The tensor product of $p$-adic Hilbert spaces

### v1

@@ 4.2 Construction of the isomorphism
    $J_{\Phi}v=v$;
@@
    $\lvert w\rangle\langle v\rvert=\mathcal{I}^{*}((J_{\Phi}v)\otimes w)$;
  3. (iii)
    $\mathcal{I}^{*}(v\otimes w)=|w\rangle\langle J_{\Phi}v|$.
@@
We observe that the relation between the $p$-adic Hilbert-Schmidt-class operators $\mathcal{T}(\mathcal{H},\mathcal{K})$ and the tensor product of $p$-adic Hilbert spaces $\mathcal{H}\otimes\mathcal{K}$ is essentially related to the correspondence $\lvert\psi\rangle\langle\phi\rvert\leftrightarrow\phi\otimes\psi,$ with necessary conjugation operators, as it happens in the complex setting.
@@ 5 Tensor product of subspaces
In what follows, for the sake of notational convenience, we set $c_{0}(I,\mathbb{Q}_{p,\mu})\equiv c_{0}(I)$ — where $I$ is a finite set, or $I=\mathbb{N}$. We start by proving
@@
where $\lambda=\{\lambda_{i}\}_{i\in I}\in c_{0}(I)$, and $x\in X$. Since $\lambda\in c_{0}(I)$, then, for every $x$ in $X$, it is clear that $\{\lambda_{i}x\}_{i\in I}$ is an element of $c_{0}(I,X)$, i.e.,
@@
We now aim to show that $\mathcal{L}_{X}$ defines a norm-preserving map between $c_{0}(I)\otimes_{\pi}X$ and $c_{0}(I,X)$. Indeed, let $u=\sum_{j=1}^{m}\lambda_{j}\otimes x_{j}$ be a possible representation of $u\in c_{0}(I)\otimes_{\pi}X$ where, without loss of generality, we can assume $\{x_{j}\}_{j=1}^{m}$ to be a norm-orthogonal system in $X$. Taking into account that, for $j=1,\ldots,m$, $\lambda_{j}=\{\lambda_{i,j}\}_{i\in I}$, we have:
@@
To prove the reverse inequality, let us fix a representation $\sum_{j=1}^{m}\lambda_{j}\otimes x_{j}$ of $u$ in $c_{0}(I)\otimes_{\pi}X$ — where, again, we assume $\{x_{j}\}_{j=1}^{m}$ to be a norm-orthogonal system in $X$. Let us consider the series $\sum_{n\in I}e_{n}\otimes u_{n}$, where $\{u_{n}\}_{n\in I}=\{\sum_{j=1}^{m}\lambda_{n,j}x_{j}\}_{n\in I}=\mathcal{L}(u)$, and where $\bm{e}=\{e_{n}\}_{n\in I}$ is the so-called canonical basis in $c_{0}(I)$ (cf. Example 2.11). Let $\Pi_{k}\colon c_{0}(I)\rightarrow c_{0}(K)$ — where $K=\{1,\ldots,k\}$, $\mathrm{card}(K)\leq\mathrm{card}(I)$ — denote the projection of the elements in $c_{0}(I)$ onto the first $k$ coordinates. We then have that
@@
It is clear that $\Pi_{k}(\lambda)\rightarrow\lambda$ for $k\rightarrow\infty$ (if $I$ is a finite set, for $k$ equal to $\text{card}(I)$), and, therefore,
@@
Thus, for $k\rightarrow\infty$ (or, if $I$ is a finite set, for $k$ equal to $\text{card}(I)$), $\|u-\sum_{n=1}^{k}e_{n}\otimes u_{n}\|_{\pi}=0$, and, hence, we proved that $\sum_{n=1}^{k}e_{n}\otimes u_{n}$ converges to $u$ in the projective norm $\|\cdot\|_{\pi}$. Now, we have:
@@
Therefore, taking into account (5) and (135), we conclude that $\|\mathcal{L}_{X}(u)\|_{\infty}=\|u\|_{\pi}$, and, hence, that $\mathcal{L}_{X}$ is a norm-preserving map (an isometry) between the $p$-adic normed spaces $c_{0}(I)\otimes_{\pi}X$ and $c_{0}(I,X)$. On the other hand, since $c_{0}(I,X)$ is complete, $\mathcal{L}_{X}$ admits a unique bounded (Hahn-Banach) extension to an isometric operator — which we still denote with the same symbol — from the completed projective tensor product $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ into $c_{0}(I,X)$.
To prove that $\mathcal{L}_{X}$ is an isomorphism of $p$-adic Banach spaces, we are left to show that it is a surjective map. To this end, let $x=\{x_{i}\}_{i\in I}$ be a sequence in $c_{0}(I,X)$. If we consider the series $\sum_{i\in I}e_{i}\otimes x_{i}$, it is clear that its image under $\mathcal{L}_{X}$ coincides with $x=\{x_{i}\}_{i\in I}$. Hence, we have only to show that $\sum_{i\in I}e_{i}\otimes x_{i}$ converges in $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$. If $\mathrm{card}(I)<\infty$, this is certainly true. On the other hand, if $I=\mathbb{N}$, since $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ is complete, it suffices to show that $\big\{\sum_{i=1}^{n}e_{i}\otimes x_{i}\big\}_{n\in\mathbb{N}}$ is a $p$-adic Cauchy sequence. This simply follows by observing that [^1] We recall that, if $(X,\|\cdot\|)$ is a $p$-adic normed space, a sequence $\{x_{n}\}_{n\in\mathbb{N}}$ is a Cauchy sequence in $X$ iff $\lim_{n}\|x_{n+1}-x_{n}\|=0$. See Chapter $1$ in [30].
@@
and recalling that, since $\{x_{n}\}_{n\in I}\in c_{0}(I,X)$, then $\lim_{n}x_{n}=0$. Hence, $\mathcal{L}_{X}$ is a surjective isometry — i.e., an isometric isomorphism — between the $p$-adic Banach spaces $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ and $c_{0}(I,X)$. ∎
@@
Let $\mathcal{H}$ be a $p$-adic Hilbert space, and let $\mathcal{W}\subset\mathcal{H}$ be a norm-closed subspace of $\mathcal{H}$. We say that $\mathcal{W}$ is a Hilbert subspace of $\mathcal{H}$ if it is a $p$-adic Hilbert space itself — equivalently, if $\mathcal{W}$ admits an orthonormal basis $\Psi\equiv\{\psi_{i}\}_{i\in I}\subset\mathcal{W}$, $\mathrm{card}(I)=\dim\mathcal{W}$. We say that $\mathcal{W}$ is a regular subspace of $\mathcal{H}$, if it is a Hilbert subspace, and if some (hence, any) orthonormal basis of $\mathcal{W}$ can be extended to an orthonormal basis of $\mathcal{H}$.
@@
$\|\mathscr{S}_{\Psi}(u)\|_{\pi}=\bigg\|\sum_{j=1}^{m}W_{\Psi}(x_{j})\otimes y_{j}\bigg\|$ $=\inf\bigg\{\max_{1\leq j\leq m}\|W_{\Psi}(x_{j})\|\,\|y_{j}\|\,\,\big\lvert\,\,u=\sum_{j=1}^{m}x_{j}\otimes y_{j},\,\,x_{j}\in\mathcal{H},y_{j}\in\mathcal{K}\bigg\}$
@@ 6 Conclusions
In this work we have introduced the tensor product of $p$-adic Hilbert spaces as a fundamental step in advancing the abstract approach theory of quantum mechanics on a quadratic extension $\mathbb{Q}_{p,\mu}$ of $p$-adic numbers [24, 22, 23]. Let us briefly summarize our main achievements:
  • We defined the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ between $p$-adic Hilbert spaces considering them as simple vector spaces, i.e. without additional structure. Then, the key point was the introduction of the norm $\|\cdot\|_{\pi},$ which is the most natural adaptation, to our case of study, of the projective norm used in tensor product of complex Banach spaces. Metrically completing the $p$-adic normed space obtained (after providing a suitable inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$ together with an orthonormal basis $\bm{e}$), we finally obtained the tensor product of $p$-adic Hilbert spaces $\Big(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi},\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi},\bm{e}\equiv\{e_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}\Big)$.
  • We characterized the space of Hilbert-Schmidt-class operators $\mathcal{T}(\mathcal{H},\mathcal{K})$ between any two $p$-adic Hilbert spaces which we know is equivalent to the space of $p$-adic trace class operators between $\mathcal{H}$ and $\mathcal{K}$.
  • We defined the conjugation operator $J_{\Phi}$ and then the set of anti-unitary operators $\overline{\mathcal{U}}(\mathcal{H})$ in a $p$-adic Hilbert space $\mathcal{H}$ as conjugations of unitary operators $\mathcal{U}(\mathcal{H})$. We used them to construct the isomorphism, similar to the one in the complex case, between the tensor product $\mathcal{H}\otimes\mathcal{K}$ and the space of trace class operators $\mathcal{T}(\mathcal{H},\mathcal{K}).$
  • We studied the tensor product of subspaces of a $p$-adic Hilbert space. We adapted some known results in the case of complex Banach spaces to our non-Archimedean case and, at the same time, we also analyzed the substantial differences. In particular, we focused on the tensor product of the $p$-adic Banach space of zero-converging sequences $c_{0}(I)$.
We left for future work the construction of the dual $(\mathcal{H}\otimes\mathcal{K})^{\prime}$ of $\mathcal{H}\otimes\mathcal{K}$, and the study of its properties, like whether or not there exists an isomorphism between $(\mathcal{H}\otimes\mathcal{K})^{\prime}$ and $\mathcal{H}\otimes\mathcal{K}$. Similarly, we did not address the construction of the bidual of $\mathcal{H}\otimes\mathcal{K}$ and, hence, the study of the reflexivity property of $\mathcal{H}\otimes\mathcal{K}$. Nevertheless, the present work paves the way for exploring the notion of entanglement in the $p$-adic setting. This can be pursued by a thorough study of the trace-class operators in $\mathcal{H}\otimes\mathcal{K}$.
@@
Here we provide detailed proofs of the results stated in Subsection 2.2. Since these theorems are proved likewise the analogous results for the single Hilbert space setting (see, in particular, Theorems 4.24.24.44.4 and $6.38$ in [22]), and only require minor modifications, we opted to place them in this appendix, so to leave the main text to the more original material. We start from
@@
Let $\mathcal{H}$ and $\mathcal{K}$ be two $p$-adic Hilbert spaces with orthonormal bases $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$, and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ respectively. From relations (22), (23), it is clear that $\mathcal{B}(\mathcal{H},\mathcal{K})\subset(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$. To prove the reverse inclusion, and, thus, the first equality in (25), we have only to observe that if $B\equiv\mathrm{op}_{\Phi,\Psi}(B_{mn})\in(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$ is an all-over matrix operator then, for any $\zeta$ in $\mathcal{H}$, we have:
@@
where, in the first equality, we used the series expansion in (24) of the bounded operator $B$ — converging in the strong operator topology — while the last quality follows from the Parseval identity (cf. (11)). Hence, from (144), and taking into account relation (145), we can conclude that $\|B\|=\sup_{m,n}\left|B_{mn}\right|=\sup_{n}\|B\phi_{n}\|$. ∎
@@
i.e., $\{\xi_{n}^{\zeta}\}_{n\in\mathbb{N}}$ is a bounded sequence. Using the isomorphism $\mathcal{L}_{\Phi}\colon\ell^{\infty}\rightarrow\mathcal{H}^{\prime}$ (see Remark 2.14), we have that
$\langle\zeta,B\chi\rangle=\sum_{n}\sum_{m}\overline{\xi_{n}^{\zeta}}\langle\phi_{n},\chi\rangle=\big(\mathcal{L}_{\Phi}(\overline{\xi^{\zeta}})\big)(\chi).$ (148)
@@
$\big(\mathcal{L}_{\Phi}(\overline{\xi^{\zeta}})\big)(\chi)=\langle\zeta,B\chi\rangle=\langle B^{\ast}\zeta,\chi\rangle,\quad\forall\chi\in\mathcal{H};$ (149)

### v2

@@ 4.2 Construction of the isomorphism
    $\langle v,\cdot\hskip 1.13809pt\rangle w=\mathcal{I}^{*}((J_{\Phi}v)\otimes w)$;
@@
    $\mathcal{I}^{*}(v\otimes w)=\langle J_{\Phi}v,\cdot\hskip 1.13809pt\rangle w$.
Proof.
For (i), it suffices to observe that, by the second assertion in Theorem 4.14, $\mathcal{I}^{\ast}=\mathcal{I}^{-1}$. Hence, applying $\mathcal{I}^{\ast}$ to both sides of (119), we have:
$\mathcal{I}^{\ast}\mathcal{I}(\langle v,\cdot\hskip 1.13809pt\rangle w)$ $=\mathcal{I}^{\ast}\big((J_{\Phi}v)\otimes w\big)$
$\langle v,\cdot\hskip 1.13809pt\rangle w$ $=\mathcal{I}^{\ast}\big((J_{\Phi}v)\otimes w\big).$ (127)
For (ii), exploiting the identity (119) in Proposition 4.13, and recalling that $J_{\Phi}$ is an involutive operator — i.e., $J_{\Phi}^{2}=\mathrm{Id}$ — we can argue that
$\langle J_{\Phi}v,\cdot\hskip 1.13809pt\rangle w$ $=\mathcal{I}^{\ast}\big(J_{\Phi}(J_{\Phi}v)\otimes w\big)$
$=\mathcal{I}^{\ast}\big((J_{\Phi}^{2}v)\otimes w\big)$
$=\mathcal{I}^{\ast}(v\otimes w).$ (128)
@@
The construction of the isomorphism provided in Theorem 4.14 — between the tensor product Hilbert space $\mathcal{H}\otimes\mathcal{K}$ and $\mathcal{T}(\mathcal{H},\mathcal{K})$ — rests, in a crucial way, on the correspondence between the compact operator $\langle v,\cdot\hskip 1.13809pt\rangle w\in\mathcal{T}(\mathcal{H},\mathcal{K})$, $v\in\mathcal{H}$, $w\in\mathcal{K}$, and the simple tensor $(J_{\Phi}v)\otimes w$. It is worth emphasizing that this isomorphism involves the conjugation operator $J_{\Phi}$ — associated with the orthonormal basis $\Phi$ in $\mathcal{H}$ — thereby mirroring the canonical antilinear isomorphism between the tensor product of complex Hilbert spaces, and the space of the Hilbert-Schmidt operators.
@@ 5 Tensor product of subspaces
In what follows, for the sake of notational convenience, we set $c_{0}(I,\mathbb{Q}_{p,\mu})\equiv c_{0}(I)$ — where $I$ is a finite set or $I=\mathbb{N}$. We start by proving the following
@@
where $\lambda=\{\lambda_{i}\}_{i\in I}\in c_{0}(I)$, and $x\in X$. Since $\lambda\in c_{0}(I)$ it is clear that $\{\lambda_{i}x\}_{i\in I}$ is an element of $c_{0}(I,X)$ for every $x$ in $X$, i.e.,
@@
We now aim to show that $\mathcal{L}_{X}$ defines a norm-preserving map between $c_{0}(I)\otimes_{\pi}X$ and $c_{0}(I,X)$. Indeed, let $u=\sum_{j=1}^{m}\lambda_{j}\otimes x_{j}$ be a possible representation of $u\in c_{0}(I)\otimes_{\pi}X$ where, without loss of generality, we can assume $\{x_{j}\}_{j=1}^{m}$ to be a norm-orthogonal system in $X$ (cf. Lemma $10.2.6$ in [29]). Taking into account that $\lambda_{j}=\{\lambda_{i,j}\}_{i\in I}$ for $j=1,\ldots,m$, we have:
@@
To prove the reverse inequality, let us fix a representation $\sum_{j=1}^{m}\lambda_{j}\otimes x_{j}$ of $u$ in $c_{0}(I)\otimes_{\pi}X$ — where, again, we assume $\{x_{j}\}_{j=1}^{m}$ to be a norm-orthogonal system in $X$. Let us consider the series $\sum_{n\in I}e_{n}\otimes u_{n}$ where $\{u_{n}\}_{n\in I}=\{\sum_{j=1}^{m}\lambda_{n,j}x_{j}\}_{n\in I}=\mathcal{L}(u)$ and $\bm{e}=\{e_{n}\}_{n\in I}$ is the so-called canonical basis in $c_{0}(I)$ (cf. Example 2.11). Let $\Pi_{k}\colon c_{0}(I)\rightarrow c_{0}(K)$ be the map — where $K=\{1,\ldots,k\}$, $\mathrm{card}(K)\leq\mathrm{card}(I)$ — denoting the projection of the elements in $c_{0}(I)$ onto the first $k$ coordinates. We then have that
@@
It is clear that $\Pi_{k}(\lambda)\rightarrow\lambda$ for $k\rightarrow\infty$ (if $I$ is a finite set, for $k$ equal to $\text{card}(I)$) and therefore
@@
Thus, for $k\rightarrow\infty$ (or, if $I$ is a finite set, for $k$ equal to $\text{card}(I)$), $\|u-\sum_{n=1}^{k}e_{n}\otimes u_{n}\|_{\pi}=0$ and hence we proved that $\sum_{n=1}^{k}e_{n}\otimes u_{n}$ converges to $u$ in the projective norm $\|\cdot\|_{\pi}$. Now, we have:
@@
Therefore, taking into account (5) and (135), we conclude that $\|\mathcal{L}_{X}(u)\|_{\infty}=\|u\|_{\pi}$, i.e., $\mathcal{L}_{X}$ is a norm-preserving map (an isometry) between the $p$-adic normed spaces $c_{0}(I)\otimes_{\pi}X$ and $c_{0}(I,X)$. On the other hand, since $c_{0}(I,X)$ is complete, $\mathcal{L}_{X}$ admits a unique bounded (Hahn-Banach) extension to an isometric operator — which we still denote with the same symbol — from the completed projective tensor product $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ into $c_{0}(I,X)$.
To prove that $\mathcal{L}_{X}$ is an isomorphism of $p$-adic Banach spaces, we are left to show that it is a surjective map. To this end, let $x=\{x_{i}\}_{i\in I}$ be a sequence in $c_{0}(I,X)$. If we consider the series $\sum_{i\in I}e_{i}\otimes x_{i}$, it is clear that its image under $\mathcal{L}_{X}$ coincides with $x=\{x_{i}\}_{i\in I}$. Hence, we have only to show that $\sum_{i\in I}e_{i}\otimes x_{i}$ converges in $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$. If $\mathrm{card}(I)<\infty$, this is certainly true. On the other hand, if $I=\mathbb{N}$, since $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ is complete, it suffices to show that $\big\{\sum_{i=1}^{n}e_{i}\otimes x_{i}\big\}_{n\in\mathbb{N}}$ is a $p$-adic Cauchy sequence. This simply follows by observing that
@@
and recalling that, since $\{x_{n}\}_{n\in I}\in c_{0}(I,X)$, $\lim_{n}x_{n}=0$. Hence, $\mathcal{L}_{X}$ is a surjective isometry — i.e., an isometric isomorphism — between the $p$-adic Banach spaces $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ and $c_{0}(I,X)$. ∎
@@
Let $\mathcal{H}$ be a $p$-adic Hilbert space, and let $\mathcal{W}\subset\mathcal{H}$ be a norm-closed subspace of $\mathcal{H}$. We say that $\mathcal{W}$ is a Hilbert subspace of $\mathcal{H}$ if it is a $p$-adic Hilbert space itself — equivalently, if $\mathcal{W}$ admits an orthonormal basis $\Psi\equiv\{\psi_{i}\}_{i\in I}\subset\mathcal{W}$, $\mathrm{card}(I)=\dim\mathcal{W}$. We say that $\mathcal{W}$ is a regular subspace of $\mathcal{H}$, if it is a Hilbert subspace, and if some (therefore, any) orthonormal basis of $\mathcal{W}$ can be extended to an orthonormal basis of $\mathcal{H}$.
@@
$\|\mathscr{S}_{\Psi}(u)\|_{\pi}$ $=\bigg\|\sum_{j=1}^{m}W_{\Psi}(x_{j})\otimes y_{j}\bigg\|$
$=\inf\bigg\{\max_{1\leq j\leq m}\|W_{\Psi}(x_{j})\|\,\|y_{j}\|\,\,\big\lvert\,\,u=\sum_{j=1}^{m}x_{j}\otimes y_{j},\,\,x_{j}\in\mathcal{H},y_{j}\in\mathcal{K}\bigg\}$
@@ 6 Conclusions
This paper is a natural continuation of our previous investigations on $p$-adic quantum mechanics [24, 22, 23, 42]. In particular, in [24, 22, 23], we introduced a model of quantum mechanics where the carrier vector space of physical states is an abstract Hilbert space over a quadratic extension $\mathbb{Q}_{p,\mu}$ of $\mathbb{Q}_{p}$. When dealing with composite systems, it is natural to conjecture that — analogously to what happens in the usual (complex) formulation of quantum mechanics — the fundamental mathematical framework should involve the tensor product of two Hilbert spaces. It is precisely this issue that motivates us to introduce, in this work, a notion of tensor product tailored to our specific $p$-adic setting.
Let us briefly summarize our main achievements:
  • We have first defined the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ of two $p$-adic Hilbert spaces — which involves solely the linear structure of these spaces. Next, we have endowed $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ with a suitable non-Archimedean norm, thereby achieving a $p$-adic normed space $\mathcal{H}\otimes_{\pi}\mathcal{K}$ out of the initial algebraic tensor product. Here, we have noted that this norm may be regarded as the ultrametric analogue of the standard projective norm associated with the algebraic tensor product of two real or complex Banach spaces — and, accordingly, we denoted the associated Banach space completion by $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. As a last step of our construction, we have endowed $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$ with the inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$, and formed a new orthonormal basis $\Xi\equiv\{\xi_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}$ out of the orthonormal bases $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ in the initial Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. Eventually, this has led us to the ($p$-adic Hilbert space) tensor product $\mathcal{H}\otimes\mathcal{K}\equiv\big(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi},\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi},\bm{\Xi}\equiv\{\xi_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}\big)$.
  • As a next step, we have characterized the set of anti-unitary operators $\overline{\mathcal{U}}(\mathcal{H})$ in a $p$-adic Hilbert space $\mathcal{H}$. To this end, we have first introduced a suitable notion of a conjugation operator $J_{\Phi}$ — associated with any orthonormal basis $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$ — defined by requiring that $J_{\Phi}$ is precisely that conjugate-linear map which leaves invariant the subset $\Phi\subset\mathcal{H}$, i.e., such that $J_{\Phi}\phi_{n}=\phi_{n}$, $\forall n\in\mathbb{N}$. An anti-unitary operator has then been defined as the composition of a unitary operator in $\mathcal{H}$, and the conjugation operator $J_{\Phi}$. After having characterized the set $\overline{\mathcal{U}}(\mathcal{H})$, we have highlighted a key difference between $p$-adic and complex conjugation operators. In fact, whereas, in the complex setting, a conjugation operator in a Hilbert space $\mathscr{H}$ can be (equivalently) defined as an involutive anti-unitary operator $Z\colon\mathscr{H}\rightarrow\mathscr{H}$ — in which case one can prove there always exits an orthonormal basis in $\mathscr{H}$ preserved by the action of $Z$ — in the $p$-adic setting, it is not always possible to construct an invariant orthonormal basis $\Phi$ in $\mathcal{H}$, such that $Z\equiv J_{\Phi}$ is the associated conjugation operator.
  • As a ‘consistency check’ of our previous construction of the $p$-adic tensor product Hilbert space $\mathcal{H}\otimes\mathcal{K}$, we have exploited the conjugation operator $J_{\Phi}$ to define a conjugate-linear isomorphism between $\mathcal{H}\otimes\mathcal{K}$, and the $p$-adic Hilbert space $\mathcal{T}(\mathcal{H},\mathcal{K})$ of trace/Hilbert-Schmidt class operators from $\mathcal{H}$ to $\mathcal{K}$. This establishes a clear analogy with the familiar tensor product structure of complex Hilbert spaces.
  • Lastly, we have studied how the tensor product $\mathcal{H}\otimes\mathcal{K}$ behaves w.r.t. the subspaces of the original $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. Here, we have first proved that, similarly to the complex setting — where one can show there exists a canonical isomorphism between a complex or real tensor product Banach space $\mathscr{X}\operatorname{\widehat{\otimes}_{\pi}}\ell_{1}$ and the sequence space $\ell_{1}(\mathscr{X})$ — given a $p$-adic Banach space $(X,\|\cdot\|)$, it is possible to construct the isomorphism $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X\cong c_{0}(I,X)$. However, unlike the complex framework, we have proved here a much stronger result: We showed that the $p$-adic tensor product operation preservers the subspace structure of a $p$-adic Hilbert space. Namely, that if $\mathcal{H}$ and $\mathcal{K}$ are two $p$-adic Hilbert spaces, and if $\mathcal{W}\subset\mathcal{K}$ is an Hilbert (or a regular) subspace of $\mathcal{K}$, then $\mathcal{H}\otimes\mathcal{W}$ is an Hilbert (resp. regular) subspace of $\mathcal{H}\otimes\mathcal{K}$.
Let us now provide a “bird’s-eye view” of some possible continuations of this work. Indeed, a first interesting question one can address at this point, is what reflexivity and dual properties characterize the tensor product $\mathcal{H}\otimes\mathcal{K}$. In fact, we already know that the dual of a $p$-adic Hilbert space $\mathcal{H}$ is not a $p$-adic Hilbert space itself — if fact, it is a $p$-adic Banach space — and we observed that $\mathcal{H}$ is, in general, not reflexive, but only pseudo-reflexive [22, 23]. Hence, a detailed study of the dual and the bidual of $\mathcal{H}\otimes\mathcal{K}$ could unveil new peculiarities of $p$-adic Hilbert spaces not observed in the standard complex setting.
Once the $p$-adic tensor product Hilbert space $\mathcal{H}\otimes\mathcal{K}$ is constructed, a further natural step would be to investigate the tensor product structure of the set of adjointable, and trace/Hilbert-Schmidt class operators in $\mathcal{H}\otimes\mathcal{K}$. In particular, the former will provide a suitable notion of an observable, and the latter, that of a state, of a composite quantum system. This will also allow us to introduce a natural notion of entangled state in the $p$-adic setting and, more in general, to investigate the separability properties of $p$-adic quantum systems.
@@
In this appendix, we present detailed proofs of the theorems stated in Subsection 2.2. Since their arguments closely mirror those of the single Hilbert space setting (see, in particular, Theorems 4.24.24.44.4 and $6.38$ in [22]), and only require minor modifications, we opted to include them here, leaving the main text to the more original material. We start from
@@
Let $\mathcal{H}$ and $\mathcal{K}$ be $p$-adic Hilbert spaces with orthonormal bases $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$, and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$, respectively. From relations (23), (24), it is clear that $\mathcal{B}(\mathcal{H},\mathcal{K})\subset(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$. To prove the reverse inclusion, and, thus, the first equality in (26), it suffices to observe that if $B\equiv\mathrm{op}_{\Phi,\Psi}(B_{mn})\in(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$ then, for any $\zeta$ in $\mathcal{H}$, we have:
@@
where, in the first equality, we used the first series expansion in (25) of the bounded operator $B$ — converging in the strong operator topology — while the last quality follows from the Parseval identity (cf. (12)). Hence, from (144), and taking into account relation (145), we can conclude that $\|B\|=\sup_{m,n}\left|B_{mn}\right|=\sup_{n}\|B\phi_{n}\|$. ∎
@@
i.e., $\{\xi_{n}^{\zeta}\}_{n\in\mathbb{N}}$ is a bounded sequence. Using the isomorphism $\mathcal{E}_{\Phi}\colon\ell^{\infty}\rightarrow\mathcal{H}^{\prime}$ (see Remark 2.14), we have that
$\langle\zeta,B\chi\rangle=\sum_{n}\sum_{m}\overline{\xi_{n}^{\zeta}}\langle\phi_{n},\chi\rangle=\big(\mathcal{E}_{\Phi}(\overline{\xi^{\zeta}})\big)(\chi).$ (148)
@@
$\big(\mathcal{E}_{\Phi}(\overline{\xi^{\zeta}})\big)(\chi)=\langle\zeta,B\chi\rangle=\langle B^{\ast}\zeta,\chi\rangle,\quad\forall\chi\in\mathcal{H};$ (149)